If y=-12 when x=9 , find y when x=-4
Answer by fcabanski(1391)

There are multiple possible answers. For example:


9 - 21 = -12. That gives -4 -21 = -25.


Another answer is 9 * -12/9 = -12. -4 * -12/9 = 48/9.